Operations Report – May 28th

SOL: May 28, 2018

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System

Notes on non-nominal systems: All three solar charge controllers now nominal. Generator was on overnight and SOC was 100% this am. 41% at 6 pm. It is as if the solar is NOT going to the batteries and if it isn’t then the power we are getting is from the generator only. I’m not sure still what is going on so the troubleshooting continues. In the meantime, we are not running any AC (as in air conditioning) systems.

Generator (hours run): Turned on at 8 pm last night and off at 8 am this morning

Solar—100 % SOC at 8 am, 41% at 6 pm

Diesel – 55 %

Propane – 57 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons

Water (trailer) – 50 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: No

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

Deimos rover used: NO

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Sojourner rover used: Yes

Hours:

Beginning charge:

Ending charge:

Currently charging: will be later tonight

Spirit rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Opportunity rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

HabCar used and why, where? Yes, for water

General notes and comments: Trash taken to town today. We also got water and accidently let the pump run after the tank was empty. So we changed pumps to the other new one I purchased when the old one broke and contaminated the water supply. I was not about to have to clean everything at this time!

We removed all the old wood support frame structure from underneath the Hab today in preparation for the all new drainpipe which will be installed tomorrow. The plan was to reskirt the Hab once this was done but I think it looks nice and I would like to keep it open to the fresh air all summer so I will consult with our contractor tomorrow and maybe we will put it off until Fall. We know have plenty of wood for a nice bonfire as a bonus.

The railing on the porch was repaired and MOOP was picked up from all over campus.

Summary of internet: Nothing to report

Summary of suits and radios: Radios were located in the box under the stairs. Only 11 of the new radios remain in operations along with 6 new ones for a total of 17 radios. However we purchased 40 new radios this year so we are missing more than half of them, probably due to them being discarded by crews who determined that they were no longer working. We will have to do a little better next year on taking care of them.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Yesterday I thought it was the 26thand it was the 27th. Can someone change the date on the report and if not, let this serve as clarification of my mistake. Time flies when you are having fun on Mars!